The business plan for KnowIt, LLC, a start-up company that manufactures portable multi-gas detectors, showed equivalent annual cash flows of $400,000 for the first 5 years. If the cash flow in year 1 was $305,000 and the constant increase thereafter was $50,000 per year, what interest rate was used in the calculation?
Answer #1

|_ A B C D E - NM + 95000 Equivalent annual Incremantal cash flows Cash flow Difference 400000 305000 400000 355000 45000 400

EXCEL FORMULA:

Year Equivalent annual cash flows 400000 400000 400000 400000 400000 Incremantal Cash flow 305000 =D3+50000 =D4+50000 =D5+500